OpenStudy (anonymous):

Which ratio is equivalent to 12 : 32?

OpenStudy (idkwut):

12: 32 can be seen as a fraction. Divide both sides by 4 and you get 3/8

OpenStudy (idkwut):

Or 3:8

OpenStudy (lauluisa):

6/16 = 12/32

OpenStudy (idkwut):

Unless you were going to give us options... so long as the decimal is equal. 12/32 = 0.375. 3/8 = 0.375, etc.

OpenStudy (lauluisa):

6/16 = 6*2/16*2 6/16 = 12/32 3/8 = 3*4/8*4 3/8 = 12/32 so, 12/32 = 6/16 = 3/8

OpenStudy (anonymous):

A.48/128 B.28/63 C.7/15 D.48/64

OpenStudy (idkwut):

Which ratio gives you 0.375 ? You can do this! :D

OpenStudy (anonymous):

A

OpenStudy (idkwut):

Good job!

OpenStudy (anonymous):

ok one more Which ratio is not equivalent to 48/64 ?

OpenStudy (anonymous):

2/3 6/8 12/16 45/60

OpenStudy (idkwut):

48/64 = 0.75 Find the value that is equal to 0.75.

OpenStudy (anonymous):

6/8

OpenStudy (idkwut):

Yes, but there are also other values that are equal to 0.75.

OpenStudy (anonymous):

12/16 45/60

OpenStudy (idkwut):

Great! :)

OpenStudy (anonymous):

so ......?

OpenStudy (lauluisa):

Your answer is 2/3 LOL it's obvious. It's the only one that it's not equal to .75

OpenStudy (lauluisa):

@IDKwut You are a great help!

OpenStudy (anonymous):

ooops my bad i put the wrong question omg hold up

OpenStudy (anonymous):

the question is Which ratio is proportional to 5/8 ?

OpenStudy (lauluisa):

45/72

OpenStudy (lauluisa):

The common denominator is 9. Do the math -- 45/72 = 5/8 (5 x9 = 45) and (9 x 8 = 72)

OpenStudy (anonymous):

yea i got it